Spectra of completely p-hyponormal operators M. Chō, Joetsu, T. Huruya, Niigata and M. Itoh, Joetsu, Japan

AMS Subject classification: 47B20 Keywords: Hilbert space, hyponormal operator, p-hyponormal operator, invariant space Abstract: In this paper we study spectral properties of completely p-hyponormal operators on a complex Hilbert space.